# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ F.1 Removing seasonality
For any site $\boldsymbol{s}_{j}$, we combine daily observations across all days as a vector and denote it by $\boldsymbol{v}_{j}=(v_{j1},\cdots,v_{jN})^{T}$ where $N=11,315$ is the number of days between 1985/01/01 and 2015/12/31. Following Huser (2021), we remove the seasonality from the Red Sea SST daily records at a fixed $\boldsymbol{s}_{j}$ via subtracting the overall trend averaged within its neighborhood of radius $r=30$ km, and then we repeat the same procedure for every other location.
More specifically, denote the index set of all location with the neighborhood of $\boldsymbol{s}_{j}$ by $\mathcal{N}_{j}=\{i:||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||<r,\;i=1,\ldots,n_{s}\}$. To get rid of the seasonality in $\boldsymbol{v}_{j}$, we first concatenate all records in the neighborhood $\{\boldsymbol{v}_{i}:i\in\mathcal{N}_{j}\}$ to get a flattened response vector $\boldsymbol{V}_{j}$; that is, $\boldsymbol{V}_{j}=(\boldsymbol{v}_{i_{1}}^{T},\boldsymbol{v}_{i_{2}}^{T},\cdots,\boldsymbol{v}_{i_{|\mathcal{N}_{j}|}}^{T})^{T}$ where $\{i_{1},\ldots,i_{|\mathcal{N}_{j}|}\}$ include all elements of $\mathcal{N}_{j}$. Thus, the length of the vector $\boldsymbol{V}_{j}$ is $|\mathcal{N}_{j}|\times N$. Second, we construct the matrix $\boldsymbol{M}=(\boldsymbol{1}_{N},\boldsymbol{t},\boldsymbol{B}_{N\times 12})$, where $\boldsymbol{t}=(1,\ldots,N)^{T}$ is used to capture linear time trend and the columns of $\boldsymbol{B}$ are $12$ cyclic cubic spline bases defined over the continuous interval $[0,366]$ evaluated at $1,\ldots,N$ modulo 365 or 366 (i.e., the day in the corresponding year). These basis functions use equidistant knots over of $[0,366]$ that help capturing the monthly-varying features. Then, we vertically stack the matrix $\boldsymbol{M}$ for $|\mathcal{N}_{j}|$ times to build the design matrix $\boldsymbol{M}_{j}$. Through simple linear regression of $\boldsymbol{V}_{j}$ on $\boldsymbol{M}_{j}$, we get the fitted values $\hat{\boldsymbol{V}}_{j}=(\hat{\boldsymbol{v}}_{i_{1}}^{T},\hat{\boldsymbol{v}}_{i_{2}}^{T},\cdots,\hat{\boldsymbol{v}}_{i_{|\mathcal{N}_{j}|}}^{T})^{T}$.
@@
$\log\boldsymbol{\epsilon}_{j}$ $=\boldsymbol{M}^{\sigma}_{j}\times(\beta_{1},\beta_{2})^{T}.$
Hence we can estimate parameters $(\beta_{1},\beta_{2})^{T}$ via optimizing the multivariate normal density function, i.e.,
$(\hat{\beta}_{1},\hat{\beta}_{2})^{T}=\operatorname*{argmin}_{(\beta_{1},\beta_{2})^{T}}\left\{-\frac{1}{2}\log\boldsymbol{1}^{T}\boldsymbol{\epsilon}^{2}_{j}-\frac{1}{2}(\boldsymbol{V}_{j}-\hat{\boldsymbol{V}}_{j})^{T}\mathrm{diag}(\boldsymbol{\epsilon}^{-2}_{j})(\boldsymbol{V}_{j}-\hat{\boldsymbol{V}}_{j})\right\}.$
Let $\hat{\boldsymbol{\epsilon}}_{j}=\exp\{\boldsymbol{M}^{\sigma}_{j}\times(\hat{\beta}_{1},\hat{\beta}_{2})^{T}\}\equiv(\hat{\boldsymbol{e}}_{i_{1}}^{T},\hat{\boldsymbol{e}}_{i_{2}}^{T},\cdots,\hat{\boldsymbol{e}}_{i_{|\mathcal{N}_{j}|}}^{T})^{T}$. Note that in defining the neighborhood of site $\boldsymbol{s}_{j}$, we also include the $j$th site. By an abuse of notation, we denote the fitted values corresponding to the $j$th site by $\hat{\boldsymbol{v}}_{j}$ and $\hat{\boldsymbol{e}}_{j}$, which correspond to the mean trend and residual standard deviations at site $\boldsymbol{s}_{j}$, respectively. Finally, the daily records at $\boldsymbol{s}_{j}$ can be de-trended by calculating
@@ F.2 Marginal distributions of the monthly maxima
After removing seasonality by normalization (see Eq. F.1), we extract monthly maxima from $\boldsymbol{v}_{j}^{*}$ at site $\boldsymbol{s}_{j}$ and denote them as $\boldsymbol{m}_{j}=(m_{j1},\ldots,m_{jn_{t}})$, in which $n_{t}=372$ is the number of months between 1985/01/01 and 2015/12/31 and $j=1\ldots,n_{s}$. Before applying our proposed model, we need to find a distribution which fits the monthly maxima well so we can transform the data to the Pareto-like distribution shown in Eq. (10). Given prior experience in analyzing monthly maxima, we propose two candidate distributions: the generalized extreme value (GEV) distribution and the general non-central $t$ distribution. To choose between them, we choose to perform $\chi^{2}$ goodness-of-fit tests due to its flexibility in choosing the degrees of freedom as well as the size of intervals.
The $\chi^{2}$ goodness-of-fit test at a site $\boldsymbol{s}_{j}\in\mathcal{S}$ proceeds as follows. First, we calculate the equidistant cut points within the range of all monthly maxima at $\boldsymbol{s}_{j}$ to get $n_{I}$ intervals. Second, we count the number of monthly maxima falling within each interval and denote them by $O_{i}$ $(i=1,\ldots,n_{I})$. Third, we fit the GEV and $t$ distributions to the block maxima series at $\boldsymbol{s}_{j}$ to get the parameter estimates. Then the expected frequencies $E_{i}$ $(i=1,\ldots,n_{I})$ is calculated by multiplying the number of monthly maxima at each site (i.e., $n_{t}$) by the probability increment of the fitted GEV or $t$ distribution in each interval (denoted by $p_{i}$). Treating the frequencies as a multinomial distribution with $n_{t}$ trials and $n_{I}$ categories, we can derive the generalized likelihood-ratio test statistic for the null hypothesis $H_{0}$ that $(p_{1},\cdots,p_{n_{I}})^{T}$ are the true event probabilities. Specifically, under the null hypothesis $H_{0}$, Wilk’s Theorem guarantees
@@ F.3 Marginal transformation
Before applying our model to monthly maxima, certain transformations need to be done to match our marginals in Section 1.3.1. When performing the goodness-of-fit tests, we already obtained the sitewise GEV parameters: $\mu_{j}$, $\sigma_{j}$, and $\xi_{j}$ for $j=1,\ldots,n_{s}$. Since monotonic transformations of the marginal distributions do not alter the dependence structure of the data input, we define $x_{jt}=F_{jt}^{-1}\{F_{\mathrm{GEV}}(m_{jt};\mu_{j},\sigma_{j},\xi_{j})\},\;t=1,\ldots,n_{t},\;j=1,\ldots,n_{s}$, in which $F_{jt}$ is the marginal distribution function of $X_{t}(\boldsymbol{s}_{j})$ displayed in Eq. (10), the function $F_{\mathrm{GEV}}(\cdot;\mu_{j},\sigma_{j},\xi_{j})$ is the distribution function of $\mathrm{GEV}(\mu_{j},\sigma_{j},\xi_{j})$, and $m_{jt}$ is the monthly maximum at site $\boldsymbol{s}_{j}$ from $t^{\mathrm{th}}$ month. Further, we have $\boldsymbol{x}_{t}=(x_{1t},x_{2t},\cdots,x_{n_{s}t})^{T}$, $t=1,\ldots,n_{t}$, which will be treated as the response in Algorithm 2. It should be noted that $F_{jt}$ is defined with the parameters $\alpha_{t}$, $\boldsymbol{\gamma}_{t}^{T}$ and $\boldsymbol{\Omega}$. Recall that the matrix $\boldsymbol{\Omega}$, defined in Eq. (D.3), contains the basis function evaluations at all locations. After updating these parameters in each iteration of the stochastic gradient descent algorithm, we need to update the values of $\{x_{jt}:t=1,\ldots,n_{t},\;j=1,\ldots,n_{s}\}$ before continuing the next iteration.

### v4

@@ F.1 Removing seasonality
For any site $\boldsymbol{s}_{j}$, we combine daily observations across all days as a vector and denote it by $\boldsymbol{v}_{j}=(v_{j1},\cdots,v_{jN})^{\top}$ where $N=11,315$ is the number of days between 1985/01/01 and 2015/12/31. Following Huser (2021), we remove the seasonality from the Red Sea SST daily records at a fixed $\boldsymbol{s}_{j}$ via subtracting the overall trend averaged within its neighborhood of radius $r=30$ km, and then we repeat the same procedure for every other location.
More specifically, denote the index set of all location with the neighborhood of $\boldsymbol{s}_{j}$ by $\mathcal{N}_{j}=\{i:||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||<r,\;i=1,\ldots,n_{s}\}$. To get rid of the seasonality in $\boldsymbol{v}_{j}$, we first concatenate all records in the neighborhood $\{\boldsymbol{v}_{i}:i\in\mathcal{N}_{j}\}$ to get a flattened response vector $\boldsymbol{V}_{j}$; that is, $\boldsymbol{V}_{j}=(\boldsymbol{v}_{i_{1}}^{\top},\boldsymbol{v}_{i_{2}}^{\top},\cdots,\boldsymbol{v}_{i_{|\mathcal{N}_{j}|}}^{\top})^{\top}$ where $\{i_{1},\ldots,i_{|\mathcal{N}_{j}|}\}$ include all elements of $\mathcal{N}_{j}$. Thus, the length of the vector $\boldsymbol{V}_{j}$ is $|\mathcal{N}_{j}|\times N$. Second, we construct the matrix $\boldsymbol{M}=(\boldsymbol{1}_{N},\boldsymbol{t},\boldsymbol{B}_{N\times 12})$, where $\boldsymbol{t}=(1,\ldots,N)^{\top}$ is used to capture linear time trend and the columns of $\boldsymbol{B}$ are $12$ cyclic cubic spline bases defined over the continuous interval $[0,366]$ evaluated at $1,\ldots,N$ modulo 365 or 366 (i.e., the day in the corresponding year). These basis functions use equidistant knots over of $[0,366]$ that help capturing the monthly-varying features. Then, we vertically stack the matrix $\boldsymbol{M}$ for $|\mathcal{N}_{j}|$ times to build the design matrix $\boldsymbol{M}_{j}$. Through simple linear regression of $\boldsymbol{V}_{j}$ on $\boldsymbol{M}_{j}$, we get the fitted values $\hat{\boldsymbol{V}}_{j}=(\hat{\boldsymbol{v}}_{i_{1}}^{\top},\hat{\boldsymbol{v}}_{i_{2}}^{\top},\cdots,\hat{\boldsymbol{v}}_{i_{|\mathcal{N}_{j}|}}^{\top})^{\top}$.
@@
$\log\boldsymbol{\epsilon}_{j}$ $=\boldsymbol{M}^{\sigma}_{j}\times(\beta_{1},\beta_{2})^{\top}.$
Hence we can estimate parameters $(\beta_{1},\beta_{2})^{\top}$ via optimizing the multivariate normal density function, i.e.,
$(\hat{\beta}_{1},\hat{\beta}_{2})^{\top}=\operatorname*{argmin}_{(\beta_{1},\beta_{2})^{\top}}\left\{-\frac{1}{2}\log\boldsymbol{1}^{\top}\boldsymbol{\epsilon}^{2}_{j}-\frac{1}{2}(\boldsymbol{V}_{j}-\hat{\boldsymbol{V}}_{j})^{\top}\mathrm{diag}(\boldsymbol{\epsilon}^{-2}_{j})(\boldsymbol{V}_{j}-\hat{\boldsymbol{V}}_{j})\right\}.$
Let $\hat{\boldsymbol{\epsilon}}_{j}=\exp\{\boldsymbol{M}^{\sigma}_{j}\times(\hat{\beta}_{1},\hat{\beta}_{2})^{\top}\}\equiv(\hat{\boldsymbol{e}}_{i_{1}}^{\top},\hat{\boldsymbol{e}}_{i_{2}}^{\top},\cdots,\hat{\boldsymbol{e}}_{i_{|\mathcal{N}_{j}|}}^{\top})^{\top}$. Note that in defining the neighborhood of site $\boldsymbol{s}_{j}$, we also include the $j$th site. By an abuse of notation, we denote the fitted values corresponding to the $j$th site by $\hat{\boldsymbol{v}}_{j}$ and $\hat{\boldsymbol{e}}_{j}$, which correspond to the mean trend and residual standard deviations at site $\boldsymbol{s}_{j}$, respectively. Finally, the daily records at $\boldsymbol{s}_{j}$ can be de-trended by calculating
@@ F.2 Marginal distributions of the monthly maxima
After removing seasonality by normalization (see Eq. (F.1)), we extract monthly maxima from $\boldsymbol{v}_{j}^{*}$ at site $\boldsymbol{s}_{j}$ and denote them as $\boldsymbol{m}_{j}=(m_{j1},\ldots,m_{jn_{t}})^{\top}$, in which $n_{t}=372$ is the number of months between 1985/01/01 and 2015/12/31 and $j=1\ldots,n_{s}$. Before applying our proposed model, we need to find a distribution which fits the monthly maxima well so we can transform the data to the Pareto-like distribution shown in Eq. (6) of the main paper. Given prior experience in analyzing monthly maxima, we propose two candidate distributions: the generalized extreme value (GEV) distribution and the general non-central $t$ distribution. To choose between them, we choose to perform $\chi^{2}$ goodness-of-fit tests due to its flexibility in choosing the degrees of freedom as well as the size of intervals.
The $\chi^{2}$ goodness-of-fit test at a site $\boldsymbol{s}_{j}\in\mathcal{S}$ proceeds as follows. First, we calculate the equidistant cut points within the range of all monthly maxima at $\boldsymbol{s}_{j}$ to get $n_{I}$ intervals. Second, we count the number of monthly maxima falling within each interval and denote them by $O_{i}$ $(i=1,\ldots,n_{I})$. Third, we fit the GEV and $t$ distributions to the block maxima series at $\boldsymbol{s}_{j}$ to get the parameter estimates. Then the expected frequencies $E_{i}$ $(i=1,\ldots,n_{I})$ is calculated by multiplying the number of monthly maxima at each site (i.e., $n_{t}$) by the probability increment of the fitted GEV or $t$ distribution in each interval (denoted by $p_{i}$). Treating the frequencies as a multinomial distribution with $n_{t}$ trials and $n_{I}$ categories, we can derive the generalized likelihood-ratio test statistic for the null hypothesis $H_{0}$ that $(p_{1},\cdots,p_{n_{I}})^{\top}$ are the true event probabilities. Specifically, under the null hypothesis $H_{0}$, Wilk’s Theorem guarantees
@@ F.3 Marginal transformation
Before applying our model to monthly maxima, certain transformations need to be done to match our marginals in Section 3.1.1. When performing the goodness-of-fit tests, we already obtained the sitewise GEV parameters: $\mu_{j}$, $\sigma_{j}$, and $\xi_{j}$ for $j=1,\ldots,n_{s}$. Since monotonic transformations of the marginal distributions do not alter the dependence structure of the data input, we define $x_{jt}=F_{jt}^{-1}\{F_{\mathrm{GEV}}(m_{jt};\mu_{j},\sigma_{j},\xi_{j})\},\;t=1,\ldots,n_{t},\;j=1,\ldots,n_{s}$, in which $F_{jt}$ is the marginal distribution function of $X_{t}(\boldsymbol{s}_{j})$ displayed in Eq. (6) of the main paper, the function $F_{\mathrm{GEV}}(\cdot;\mu_{j},\sigma_{j},\xi_{j})$ is the distribution function of $\mathrm{GEV}(\mu_{j},\sigma_{j},\xi_{j})$, and $m_{jt}$ is the monthly maximum at site $\boldsymbol{s}_{j}$ from $t^{\mathrm{th}}$ month. Further, we have $\boldsymbol{x}_{t}=(x_{1t},x_{2t},\cdots,x_{n_{s}t})^{\top}$, $t=1,\ldots,n_{t}$, which will be treated as the response in Algorithm 2. It should be noted that $F_{jt}$ is defined with the parameters $\alpha_{t}$, $\boldsymbol{\gamma}_{t}$ and $\boldsymbol{\Omega}$. Recall that the matrix $\boldsymbol{\Omega}$, defined in Eq. (D.3), contains the basis function evaluations at all locations. After updating these parameters in each iteration of the stochastic gradient descent algorithm, we need to update the values of $\{x_{jt}:t=1,\ldots,n_{t},\;j=1,\ldots,n_{s}\}$ before continuing the next iteration.
@@ F.4 Empirical $\chi_{h}(u)$ estimates
F.5 Additional results
Figure F.3 shows emulated replicates of the original monthly maxima field for the first and last months (1985/01 and 2015/12, respectively). Here, we convert the emulated values back to the original data scale using the estimated GEV parameters fitted from the previous step. Figure F.3 demonstrates that the XVAE is able to capture the detailed features of the temperature fields and to accurately characterize spatial dependence, while the QQ-plot shows an almost perfect alignment with the 1-1 line.
Refer to caption Refer to caption Refer to caption
Figure F.3: Observed (left) and emulated (middle) Red Sea SST monthly maxima, for the 1985/01 (top) and 2015/12 (bottom) months. From the emulation maps and QQ plots (right), we see that the emulated fields from the XVAE match the observations very well.